OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (–3, 2) and has a slope of –2? y = __ x + __

OpenStudy (anonymous):

do you know the slope intercept form of equation of a line ?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

any straight line can be represented in the form of \[y = mx + c\] where 'm' is called the slope of the line and 'c' is called the intercept.

OpenStudy (anonymous):

Can you try now ?

OpenStudy (anonymous):

would it be y=-2x+c to start with?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

find the value of 'c' by substituting the point and you have your equation :)

OpenStudy (anonymous):

would it be y=-2x+2?? sorry i'm really bad at these :/

OpenStudy (anonymous):

@xavier123

OpenStudy (anonymous):

y = -2x+c for (-3,2) 2 = -2(-3)+c => 2 = 6 + c => c = -4 hence, y= -2x-4

OpenStudy (anonymous):

oh okay... Thank you!

OpenStudy (anonymous):

most welcome :)